# Feedback-driven recurrent quantum neural network universality

### v1

@@ 2.1 The recurrent quantum neural network (RQNN) construction
Our recurrent quantum circuit is constructed based on two parametric quantum gates $\mathtt{U}$ and $\mathtt{V}$, which we now introduce. The construction builds on the feedforward quantum neural network architecture introduced in [GJ25].
@@
$\begin{array}[]{cl}\mathtt{U}^{(i)}_{1}(\bm{x},\bm{z})&:=\mathtt{H}\,\mathtt{R}_{\mathrm{z}}\left(-b^{i}\right)\mathtt{R}_{\mathrm{z}}\left(-a^{i}_{N+d}z_{d}\right)\cdots\mathtt{R}_{\mathrm{z}}\left(-a^{i}_{N+1}z_{1}\right)\mathtt{R}_{\mathrm{z}}\left(-a^{i}_{N}x_{N}\right)\cdots\mathtt{R}_{\mathrm{z}}\left(-a^{i}_{1}x_{1}\right)\mathtt{H}=\mathtt{R}_{\mathrm{x}}\left(\delta^{i}\right)\end{array}$
for any $\bm{x}=(x_{1},\ldots,x_{N})\in\mathbb{R}^{N}$ and $\bm{z}=(z_{1},\ldots,z_{d})\in\mathbb{R}^{d}$, where $\mathtt{H}$ is the Hadamard gate and
$\delta^{i}:=-b^{i}-a^{i}_{N+d}z_{d}\dots-a^{i}_{N+1}z_{1}-a^{i}_{N}x_{N}\dots-a^{i}_{1}x_{1}.$
@@
Here, $n_{0}$ is chosen as the smallest natural number such that the matrix dimension $n_{\mathtt{U}}=4n+n_{0}$ is a power of $2$, that is, $n_{\mathtt{U}}=2^{\mathfrak{n}}$. It can be easily shown that $n_{0}=4\kappa$ with $\kappa\in\mathbb{N}$, since $4n+n_{0}$ and $22n+n_{0}/$ 2 must be even for $\mathfrak{n}\geq 2$. Then, $\mathtt{U}\in\mathbb{C}^{n_{\mathtt{U}}\times n_{\mathtt{U}}}$ is a unitary quantum gate operating on $\mathfrak{n}=\log_{2}(n_{\mathtt{U}})=2+\log_{2}(n+\kappa)$ qubits with a diagonal-block structure:
$\mathtt{U}_{\bm{\theta}}(\bm{x},\bm{z})=\sum^{n-1}_{i=0}\ket{i}\bra{i}\otimes\bar{\mathtt{U}}^{(j+1)}(\bm{x},\bm{z})+\sum^{n+\kappa-1}_{i=n}\ket{i}\bra{i}\otimes{\bf 1}_{4\times 4}.$
These unitary operators with a block structure are known as uniformly controlled quantum gates. They are present in many quantum algorithms and are used to decompose general unitary gates and locally prepare arbitrary quantum states [MVBS04a, MVBS04b, BVMS05, ADMQ+22, PPR19]. They are defined as multi-controlled unitaries where each unitary block targets a set of qubits, two qubits in this case, while the other $\log_{2}(n+\kappa)$ qubits act as control qubits. Notice that the block structure of the unitary $\mathtt{U}_{\bm{\theta}}$ arises from indexing the targets as the lowest-order bits. Recently, efficient decompositions of multi-controlled unitaries have been proposed in terms of the number of single-qubit and two-qubit gates [ZB24, ZB25], as well as for approximations of the multi-controlled gate [SAAdS24]. Code implementations of these quantum gates can be found in the Qclib library [AAdS+23]. Finally, the identity blocks ${\bf 1}_{4\times 4}$ do not introduce additional gates into the quantum circuit, so the effective circuit can be reduced to the application of the $\bar{\mathtt{U}}^{(i)}$ gates. However, the number of control qubits is fixed by $\log_{2}(n+\kappa)$ and we need all of them to compute the output probabilities, as we will see below.
@@
Next, let $\mathtt{V}\in\mathbb{C}^{n_{\mathtt{U}}\times n_{\mathtt{U}}}$ be any unitary matrix mapping |$\ket{0}^{\otimes\mathfrak{n}}$ to the state |$\ket{\psi}=\frac{1}{\sqrt{n}}\sum_{i=0}^{n-1}\ket{4i}$ which for $n\geq 2$ is explicitly given by |$\ket{\psi}=\frac{1}{\sqrt{n}}\sum_{i=0}^{n-1}\ket{i}\otimes\ket{00}$. We refer to [GJ25] for a more detailed discussion on the choice of $\mathtt{V}$. Notice that when $n_{0}=0$, there is an explicit construction of $V$ in terms of Hadamard gates acting on the control qubits, see Appendix A in [GJ25].
@@
We can now measure the state of the $\mathfrak{n}$-qubit system after applying the gates $\mathtt{V}$ and $\mathtt{U}$. The possible states that we could measure are given by ${0,\ldots,n_{\mathtt{U}}-1}$ (in binary). By running the circuit repeatedly, we can now obtain (up to well-controlled Monte Carlo error, see [GJ25]) the probabilities $\mathbb{P}_{m}^{n}$ that the measured state is in $\{m,4+m,\ldots,4(n-1)+m\}$, for $m\in\{0,1,2,3\}$, where $m$ is the binary state of the last two qubits (the target qubits).
More formally, consider the unitary gate map $\mathtt{C}(\bm{x},\bm{z})=\mathtt{C}_{\mathfrak{n},\bm{\theta}}(\bm{x},\bm{z}):=\mathtt{U}_{\bm{\theta}}(\bm{x},\bm{z})\mathtt{V}$ acting on $\mathfrak{n}=2+\log_{2}(n+\kappa)$ qubits. This circuit acts on the initial state |$\ket{0}^{\otimes\mathfrak{n}}$ via the quantum gates $\mathtt{V}$ and $\mathtt{U}$ as
$\mathtt{C}_{\mathfrak{n},\bm{\theta}}(\bm{x},\bm{z})\ket{0}^{\otimes\mathfrak{n}}=\sum^{n-1}_{i=0}\ket{i}\otimes\mathtt{U}^{(i+1)}_{1}(\bm{x},\bm{z})\ket{0}\otimes\mathtt{U}^{(i+1)}_{2}\ket{0}.$
@@
$\mathbb{P}_{m}^{n,\bm{\theta}}=\mathbb{P}_{m}^{n,\bm{\theta}}(\bm{x},\bm{z}):=\mathbb{P}\left("\mathtt{C}_{\mathfrak{n},\bm{\theta}}(\bm{x},\bm{z})\ket{0}^{\otimes\mathfrak{n}}\in\{m,4+m,\ldots,4(n-1)+m\}"\right).$ (2)
This is the sum of the probabilities of being in the states |$\ket{i}\otimes\ket{m}$, where $i=0,\dots,n-1$. That is,
$\mathbb{P}_{m}^{n,\bm{\theta}}(\bm{x},\bm{z})=\sum^{n}_{i=1}\left|\bra{m}\left(\mathtt{U}^{(i)}_{1}(\bm{x},\bm{z})\ket{0}\otimes\mathtt{U}^{(i)}_{2}\ket{0}\right)\right|^{2}.$
@@
With $n$ (or equivalently $\mathfrak{n}$) fixed, the quantum circuit introduced above is uniquely defined by the choice of circuit parameters $\bm{\theta}\in\bm{\Theta}$. In what follows, we will now run $N$ such circuits in parallel. Each circuit is described by its parameters $\bm{\theta}^{j}\in\bm{\Theta}$, $j\in\left\{1,\ldots,N\right\}$. The circuit outputs then induce maps $\mathbb{P}_{m}^{n,\bm{\theta}^{j}}\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to[0,1]$ by the circuit output probabilities (2) with parameters for the $j$-th circuit given by $\bm{\theta}=\bm{\theta}^{j}$.
@@
With these ingredients, we can now define the recurrent quantum neural network that we will consider. Given the gate map $\mathtt{C}_{\mathfrak{n},\bm{\theta}}$ and $R>0$, we define $\bar{F}^{n,\bm{\theta}}_{R}:\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}^{N}$ by its component maps $\bar{F}^{n,\bm{\theta}}_{R}=(\bar{F}^{n,\bm{\theta}}_{R,1},\ldots,\bar{F}^{n,\bm{\theta}}_{R,N})$. For $j=1,\ldots,N$, the $j$-th component map $\bar{F}^{n,\bm{\theta}}_{R,j}:\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}$ is defined by
@@ 2.2 RQNN approximation of state-space maps and their derivatives
$\mathcal{F}$ :=$:=\Big{\{}f:\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}:f\in\mathcal{C}\left(\mathbb{R}^{N}\times\mathbb{R}^{d}\right)\cap L^{1}\left(\mathbb{R}^{N}\times\mathbb{R}^{d}\right),\quad\|\widehat{f}\|_{1}<\infty\Big{\}},$ (5)
$\mathcal{F}_{R}$ :={f, $:=\left\{f\in\mathcal{F},\text{ with }\|\widehat{f}\|_{1}\leq R\right\},\qquad\text{for }R>0.$ (6)
@@
$\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})=\frac{1}{n}\sum_{i=1}^{n}R\cos\left(\gamma^{i,j}\right)\cos\left(b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})\right)$ (7)
for all $(\bm{x},\bm{z})\in\mathbb{R}^{N}\times\mathbb{R}^{d}$.
Proof.
It follows directly from Proposition 1 in [GJ25]. ∎
Let $\mu$ be an arbitrary probability measure on $(\mathbb{R}^{N}\times\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{N}\times\mathbb{R}^{d}))$. Recall the notation
@@
Let $R>0$ and suppose $F=(F_{1},\ldots,F_{N}):\mathbb{R}^{N}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ is continuously differentiable and satisfies $F_{j}\in\mathcal{F}_{R}$ and $\partial_{i}F_{j}\in\mathcal{F}$ and $\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\xi_{i}^{2}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$ for $i=1,\ldots,N+d$ and $j=1,\ldots,N$. Then, for any $n\in\mathbb{N}$, there exists $\bm{\theta}\in\bm{\Theta}^{N}$ such that for any $j\in\{1,\ldots,N\}$,

### v2

@@ 3 Recurrent quantum neural network architecture
Next, we introduce in detail the employed RQNN architecture. Our recurrent quantum circuit is constructed based on two parametric quantum gates $\mathtt{U}$ and $\mathtt{V}$, which we now introduce. The construction extends the feedforward QNN architecture introduced in Gonon & Jacquier (2025) to a recurrent setting by feeding back the network’s state.
@@
$\begin{array}[]{cl}\mathtt{U}^{(i)}_{1}(\bm{x},\bm{z})&:=\mathtt{H}\,\mathtt{R}_{\mathrm{z}}\left(-b^{i}\right)\mathtt{R}_{\mathrm{z}}\left(-a^{i}_{N+d}z_{d}\right)\cdots\mathtt{R}_{\mathrm{z}}\left(-a^{i}_{N+1}z_{1}\right)\mathtt{R}_{\mathrm{z}}\left(-a^{i}_{N}x_{N}\right)\cdots\mathtt{R}_{\mathrm{z}}\left(-a^{i}_{1}x_{1}\right)\mathtt{H}\end{array}$
for any $\bm{x}=(x_{1},\ldots,x_{N})\in\mathbb{R}^{N}$ and $\bm{z}=(z_{1},\ldots,z_{d})\in\mathbb{R}^{d}$, with $\mathtt{H}$ the Hadamard gate. We may rewrite
$\mathtt{U}^{(i)}_{1}(\bm{x},\bm{z})=\mathtt{R}_{\mathrm{x}}\left(\delta^{i}\right),\quad\delta^{i}:=-b^{i}-a^{i}_{N+d}z_{d}\dots-a^{i}_{N+1}z_{1}-a^{i}_{N}x_{N}\dots-a^{i}_{1}x_{1}.$
@@
Here, $n_{0}$ is chosen as the smallest natural number such that the matrix dimension $n_{\mathtt{U}}=4n+n_{0}$ is a power of $2$, that is, $n_{\mathtt{U}}=2^{\mathfrak{n}}$. It can be easily shown that $n_{0}=4\kappa$ with $\kappa\in\mathbb{N}$, since $4n+n_{0}$ and $2n+n_{0}/2$ must be even for $\mathfrak{n}\geq 2$. Then, $\mathtt{U}\in\mathbb{C}^{n_{\mathtt{U}}\times n_{\mathtt{U}}}$ is a unitary quantum gate operating on $\mathfrak{n}=\log_{2}(n_{\mathtt{U}})=2+\log_{2}(n+\kappa)=\lceil\log_{2}(2n)\rceil$ qubits with a diagonal-block structure:
[…]
These unitary operators with a block structure are known as uniformly controlled quantum gates. They are present in many quantum algorithms and are used to decompose general unitary gates and locally prepare arbitrary quantum states (Möttönen et al., 2004; Mottonen et al., 2004; Bergholm et al., 2005; Arrazola et al., 2022; Park et al., 2019). They are defined as multi-controlled unitaries where each unitary block targets a set of qubits, two qubits in this case, while the other $\log_{2}(n+\kappa)$ qubits act as control qubits. Multi-controlled unitaries are applied depending on the state of the control qubits, which are unchanged, and only modify the target qubits. These operations generalize the CNOT gate for two qubits, in that we can now have several control and target qubits. Notice that the block structure of the unitary $\mathtt{U}_{\bm{\theta}}$ arises from indexing the targets as the lowest-order bits. Recently, efficient decompositions of multi-controlled unitaries have been proposed in terms of the number of single-qubit and two-qubit gates (Zindorf & Bose, 2024; 2025), as well as for approximations of the multi-controlled gate (Silva et al., 2024). Code implementations of these quantum gates can be found in the Qclib library (Araujo et al., 2023). Finally, the identity blocks ${\bf 1}_{4\times 4}$ do not introduce additional gates into the quantum circuit, so the effective circuit can be reduced to the application of the $\bar{\mathtt{U}}^{(i)}$ gates. However, the number of control qubits is fixed by $\log_{2}(n+\kappa)$ and we need all of them to compute the output probabilities, as we will see below.
@@
[…]
@@
We can now measure the state of the $\mathfrak{n}$-qubit system after applying the gates $\mathtt{V}$ and $\mathtt{U}$. The possible states that we could measure are given by $0,\ldots,n_{\mathtt{U}}-1$} (in binary). By running the circuit repeatedly, we can now obtain (up to well-controlled Monte Carlo error, see Appendix E) the probabilities $\mathbb{P}_{m}^{n}$ that the measured state is in $\{m,4+m,\ldots,4(n-1)+m\}$, for $m\in\{0,1,2,3\}$, where $m$ is the binary state of the last two qubits (the target qubits).
[…]
@@
[…]
@@
With $n$ (or equivalently $\mathfrak{n}$) fixed, the quantum circuit introduced above is uniquely defined by the choice of circuit parameters $\bm{\theta}\in\bm{\Theta}$. We will now run $N$ such circuits in parallel, each representing a component of the state map $F$ in (1). Each circuit is described by its parameters $\bm{\theta}^{j}\in\bm{\Theta}$, $j\in\left\{1,\ldots,N\right\}$. The circuit outputs then induce maps $\mathbb{P}_{m}^{n,\bm{\theta}^{j}}\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to[0,1]$ by the circuit output probabilities (2) with parameters for the $j$-th circuit given by $\bm{\theta}=\bm{\theta}^{j}$.
Figure 1: Schematic representation of the $j$-th circuit given by parameters $\bm{\theta}^{j}\in\bm{\Theta}$, $j\in\left\{1,\ldots,N\right\}$.
@@
With these ingredients, we can now define the RQNN that we will consider. Given the gate map $\mathtt{C}_{\mathfrak{n},\bm{\theta}}$ and $R>0$, we define $\bar{F}^{n,\bm{\theta}}_{R}:\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}^{N}$ by its component maps $\bar{F}^{n,\bm{\theta}}_{R}=(\bar{F}^{n,\bm{\theta}}_{R,1},\ldots,\bar{F}^{n,\bm{\theta}}_{R,N})$. For $j=1,\ldots,N$, the $j$-th component map $\bar{F}^{n,\bm{\theta}}_{R,j}:\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}$ is defined by
@@
Figure 1 provides a schematic representation of how the RQNN acts at each time step for the $j$-th circuit: at any time $t$, the system is initialized, the gates $\mathtt{V}$ and $\mathtt{U}_{\bm{\theta}^{j}}(\hat{\bm{x}}_{t-1},\bm{z}_{t})$ are applied, and the system is measured. This process is repeated to estimate the probabilities $\mathbb{P}_{1}^{n,\bm{\theta}^{j}}(\hat{\bm{x}}_{t-1},\bm{z}_{t})$ and $\mathbb{P}_{2}^{n,\bm{\theta}^{j}}(\hat{\bm{x}}_{t-1},\bm{z}_{t})$, which are aggregated into the network output $\bar{F}^{n,\bm{\theta}}_{R,j}(\hat{\bm{x}}_{t-1},\bm{z}_{t})$ according to (3). Once this is done for all $j\in\{1,\dots,N\}$, the network state $\hat{\bm{x}}_{t}$ is stored to be used as feedback for the next time step $t+1$.
@@
4 Recurrent quantum neural network universality
This section contains approximation guarantees and universality results for the recurrent quantum neural network (RQNN) family. To achieve this, in Section 4.1 we first prove refined approximation error bounds (that generalize those in Gonon & Jacquier (2025)) for feedforward quantum neural networks (QNNs) that allow us to control the error committed when approximating a function and its derivatives simultaneously, a crucial ingredient for analysing the RQNN feedback loop. These error bounds show how recurrent QNNs can be used to approximate state-space maps $F$ arbitrarily well as long as these are sufficiently smooth and satisfy Barron-type integrability conditions like, for example, $\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\xi_{i}^{2}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$, for $i=1,\ldots,N+d$ and $j=1,\ldots,N$, or $I_{q}=\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\|\bm{\xi}\|^{q}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$, for some $q\geq 2$ (see Proposition 4.2 and Corollary 4.3 below); RQNN state maps are hence universal in that category. These bounds are devised with respect to $L^{\infty}$ and $L^{2}$-type norms. As we shall prove, in the $L^{\infty}$ case, the universality of the RQNN family still holds with respect to state maps that do not necessarily satisfy the Barron condition, even though in that framework we do not formulate approximation bounds. Finally, in the last two sections, we exploit all these results on the approximation of state maps to obtain universality statements and error bounds for the approximation of arbitrary causal, time-invariant, and fading memory filters using a modified recurrent QNN. In addition to the tools developed here, our proofs of these results rely on techniques from Gonon & Ortega (2020; 2021) and the overall strategy is reminiscent of the so-called internal approximation approach introduced in Grigoryeva & Ortega (2018b, Theorem 3.1 (iii)) for echo state networks, which consists of obtaining approximation results for filters out of statements of that type for the state maps that generate them.
The approximation rate in all our results is free from the curse of dimensionality: the error decays as $\frac{1}{\sqrt{n}}$ as we increase $n$, with this rate of decay being independent of the input dimension $d$ and the state space dimension $N$. Moreover, the required number of qubits $\mathfrak{n}=\lceil\log_{2}(2n)\rceil$ is only growing logarithmically in the accuracy parameter $n$. Put differently, our circuit requires only $\mathcal{O}(\varepsilon^{-2})$ weights and $\mathcal{O}(\lceil\log_{2}(\varepsilon^{-1})\rceil)$ qubits suffices to achieve approximation error $\varepsilon>0$ when approximating functions with sufficiently integrable Fourier transforms.
@@ 4.1 RQNN approximation of state-space maps and their derivatives
$\mathcal{F}$ $:=\Big\{f:\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}:f\in\mathcal{C}\left(\mathbb{R}^{N}\times\mathbb{R}^{d}\right)\cap L^{1}\left(\mathbb{R}^{N}\times\mathbb{R}^{d}\right),\quad\|\widehat{f}\|_{1}<\infty\Big\},$ (5)
$\mathcal{F}_{R}$ $:=\left\{f\in\mathcal{F},\text{ with }\|\widehat{f}\|_{1}\leq R\right\},\qquad\text{for }R>0.$ (6)
@@
Our first result derives a representation for the QRNN output.
@@
$\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})=\frac{1}{n}\sum_{i=1}^{n}R\cos\left(\gamma^{i,j}\right)\cos\left(b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})\right),\quad\text{for all}~(\bm{x},\bm{z})\in\mathbb{R}^{N}\times\mathbb{R}^{d}.$ (7)
Our next result provides an approximation error bound for the QRNN state map jointly with its derivatives. The proof is provided in Appendix B.2. Let $\mu$ be an arbitrary probability measure on $(\mathbb{R}^{N}\times\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{N}\times\mathbb{R}^{d}))$. Recall the notation
@@
Let $R>0$ and suppose $F=(F_{1},\ldots,F_{N}):\mathbb{R}^{N}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ is continuously differentiable and satisfies $F_{j}\in\mathcal{F}_{R}$ and $\partial_{i}F_{j}\in\mathcal{F}$ and $\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\xi_{i}^{2}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$ for $i=1,\ldots,N+d$ and $j=1,\ldots,N$. Then, for any $n\in\mathbb{N}$, there exists $\bm{\theta}\in\bm{\Theta}^{N}$ such that